The coordinates below represent two linear equations.

How many solutions does this system of equations have?


Line 1
x y
–4 3
–2 4

Line 2
x y
0 8
2 6








A.

0





B.

exactly 1





C.

exactly 2





D.

infinitely many


Answer by stanbon(75887)

Line 1
x y
–4 3
–2 4
---
slope = (4-3)/(-4--2) = 1/-2 = -1/2
--------------------
Line 2
x y
0 8
2 6
----
slope = (6-8)/(2-0) = -2/2 = -1
--------
Since the slopes are different, the line meet at one point.
Ans:: One solution for the system
---------
Cheers,
Stan H.
------------